# A Generation Framework with Strict Constraints for Crystal Materials Design

### v2

@@ A.2 Crystal Symmetry Constraint
Crystal symmetry, encompassing space groups and Wyckoff positions, plays a pivotal role in the description of crystal structures. We illustrates the influence of space group types on lattice parameters in Table 5. Space group types are numbered from 0 to 230 and correspond to the seven crystal systems. Among them, Space groups with higher numerical indices impose more stringent lattice constraints. For instance, 195 $\sim$ 230 belong to the cubic system, where the lattice adopts a cube-shaped configuration with only one degree of freedom in lattice parameters. Conversely, 1$\sim$ 2 correspond to the triclinic system, featuring lattices of arbitrary geometry with six degrees of freedom. The trigonal system presents a unique case, incorporating both rhombohedral and hexagonal lattice configurations, thereby encompassing two distinct types of lattice constraints.
@@
Trigonal 143 $\sim$ 167 $a=b=c$ $a=b\neq c$ $\alpha=\beta=\gamma\neq 90^{\circ}$ $\alpha=\beta=90^{\circ},\gamma=120^{\circ}$
@@
Appendix B Implementation Details
B.1 Hyper-parameters and Training Details
@@
We train CDVAE [13], DiffCSP++ [16], and our crystal structure generator on three public datasets to validate the effectiveness of the structure generator. Using the dataset described in Section §4.1, we train Con-CDVAE [18], MatterGen [17], and our CrystalGF to verify the validity of the overall framework. We split our data into 80% training, 10% validation, and 10% test sets while maintaining strict correspondence between space group and Wyckoff positions entries, this opeartion prevents data leakage by ensuring identical crystals are allocated to the same subset. To the best of our knowledge, Con-CDVAE and MatterGen represent the advanced multi-conditional generative models, which utilize chemical formulas with material properties and elemental composition with material properties as model inputs, respectively. Therefore, we conduct comparative evaluations of both models against our CrystalGF.
@@
For CDVAE, we apply the transformer-enhanced [45] version of CDVAE as introduced in DiffCSP [14], while maintaining consistent training parameters. For DiffCSP++, we train the denoising model with 6 layers, 512-dimension hidden layer, and 128-dimension Fourier embeddings and the training epochs are set to 3000, 1500, 1000 for Perov-5, MP-20, and MPTS-52. For MatterGen, we fine-tuning the pre-trained model with chemical system and property, including band gap and formation energy. The dimension of the property embedding and the chemical system multi-hot embedding is set to 512, and model is trained for 1000 epochs with an Adam optimizer with initial learning rate $5\times 10^{-6}$ and a Plateau scheduler with a decaying factor 0.6 and a patience of 100 epochs. For Con-CDVAE, we train model with chemical formula and property, including band gap and formation energy. The formula embedding module has 3 layers, 92-dimension input layer and 64-dimension hidden layer, the property embedding module has 3 layers and 64-dimension hidden layer and the model is trained for 1000 epochs with an Adam optimizer with initial learning rate $1\times 10^{-3}$ and a Plateau scheduler with a decaying factor 0.6 and a patience of 30 epochs.
@@
For our symmetry generator, We fine-tune four LLMs ,including Qwen2.5-7B-Instruct, Llama-3.1-8B-Instruct, DeepSeek-R1-8B, and GLM-4-9B-chat, using LoRA from Llamafactory [39] framework. The LoRA fine-tuning specifically targets the qkv layers in transformer modules with 20 epochs and a $1\times 10^{-4}$ initial learning rate, the learning rate scheduler type is set to cosine annealing strategy with 0.1 warm up ratio. And the fine-tuning precision is set to FP32 throughout all computations.
@@
B.2 Symmetry Accuracy Details
@@
When constructing the confusion matrix for Wyckoff positions generation, we first identify matches where both the atomic order and generated letters coincide. Second, we distinguish cases where generated letters match despite atomic order discrepancies. Finally, atoms with inconsistent generation quantities are handled by padding with an "other" category.
B.3 DFT Calculation Details
@@
CDVAE Con-CDVAE DiffCSP++ MatterGen Constraint Generator Structure Generator NVIDIA A40 GPU 12GB 24GB 24GB 45GB 26GB + 30GB 30GB 10h 12h 30h 50h 9h + 18h 46h
B.4 Experiments Compute Resources
We present the computational resources consumed for training in Table 6, including GPU type, memory usage, and training duration. The resource consumption for CDVAE and DiffCSP++ corresponds to training with the MP-20 dataset mentioned in §4.2, while Con-CDVAE, MatterGen, the Constraint Generator, and the Structure Generator reflect the resource usage when trained on the dataset described in §4.1.
@@ B.4 Experiments Compute Resources
Appendix C Visualization of Generated Crystals Structures
@@
Appendix D Societal Impacts
@@ Appendix E Limitations
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In the abstract and the latter part of the introduction, we explicitly state the scope of this work as crystal structure generation. The principal contribution of this paper lies in proposing a two-stage strict constrained generation framework, which demonstrates a 16% reduction in RMSE for generating crystals with target property compared to state-of-the-art approaches.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of the proposed method in Appendix E.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We provide a comprehensive description of our crystal generation framework in §3, comprising a LLMs-based constraint generator and a diffusion-based structure generator. In §4 ,we detail the datasets, the experimental process for obtaining results, along with required Python Class and parameter configurations. In Appendix B.1, we document the model hyper-parameters and hardware specifications used during training.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We follow the NeurIPS code and data submission guidelines, and provide the datasets used in this paper, the URL to access the model code, model weights, and inference scripts in the supplementary materials.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In §4.2 and Appendix B.1, we provide detailed descriptions of dataset splits, along with the hyper-parameters configurations and optimizer types employed during model training.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We describe the statistical methods for symmetry precision in Appendix B.2, present the calculation methods of the matching rate and RMSE in crystal structure generation in §4.2, and introduce the absolute error probability distribution and RMSE computation in property error evaluation in §4.3.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In Appendix B.4, we detail the computational resources consumed during the experiments, including compute workers, memory allocation, and execution time.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have thoroughly read and understood the NeurIPS Code of Ethics, and affirm that the research presented in this paper complies with the NeurIPS Code of Ethics in all aspects.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: In Appendix D, we discuss both the positive and negative societal impacts of the proposed crystal structure generation framework.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our data are exclusively sourced from open-source, risk-free crystal structure datasets. Furthermore, this study does not involve any high-risk scenarios with potential for model misuse, such as large language model pre-training or image generation.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We confirm that all assets used in the paper, including code, data, and models, have been properly credited, and the license and usage terms have been explicitly mentioned and complied with.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide the assets introduced in this paper in the supplementary material, including the dataset, URLs to model code, model weights, inference scripts, and corresponding documentation.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper focuses on crystal structure generation, and does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper focuses on crystal structure generation, and does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: We have read and complied with the NeurIPS LLM Policy. Specifically, we elaborate on the usage of LLMs in §3.2, detail the tools and parameters for fine-tuning LLMs in Appendix B.1, and present the testing results of the LLMs in §3.2.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Supplementary Material
In the supplementary materials, we provide the data, source code, weight files, scripts, and partial experimental results used in this work. Specifically, the CrystalGF file contains the source code, code environment, yaml files, scripts, and some experimental results of our proposed method. Pretrained model and lora weights can be downloaded at https://modelscope.cn/models/chachapro/CrystalGF. Data and code can be found at https://github.com/chachapro/CrystalGF. For Con-CDVAE and MatterGen, we provide the training results used for the experimental comparisons in this paper. Meanwhile, we have ensured that all content within the supplementary materials and website links has been anonymized.

### v3

@@ A.2 Crystal Symmetry Constraint
[…]
@@
Trigonal 143 $\sim$ 167
$a=b=c$
$a=b\neq c$
$\alpha=\beta=\gamma\neq 90^{\circ}$
$\alpha=\beta=90^{\circ},\gamma=120^{\circ}$
@@
Appendix B Our Method CrystalGF
we present CrystalGF, a novel framework for crystal materials generation that synthesizes desired crystal structures guided by chemical composition and material properties.
B.1 Architecture
A crystal generation model conditioned with chemical composition and material properties could be described as M=H(C,P){M}=H(C,P), where $H(.)$ represents a crystal structure generator, $C=({c_{1}},{c_{2}},...,{c_{k}})$ are the chemical composition constraint in which k components are specified, P is property constraint. $M=(A,X,L)$ represents for crystal structure, where $A=[{a_{1}},{a_{2}},...,{a_{n}}]$ are the elemental composition of atoms inside a unit cell, $X=[{x_{1}},{x_{2}},...,{x_{n}}]\in{{[0,1)}^{3\times n}}$ are fractional coordinates of $n$ atoms, $L=[{l_{1}},{l_{2}},{l_{2}}]\in{\mathbb{R}^{3\times 3}}$ is the lattice. In this work, we formulate our model as:
$\begin{array}[]{c}{M}=H(G(C,P),P)\end{array}$
Notice that for our model, the actual input constraints are still chemical composition and material properties, and we utilize a constraint generator to obtain additional constraints. We utilize a LLMs-based model as our constraint generation module (Appendix B.2), and a multi-conditional constrained diffusion model as the crystal structure generation module (Appendix B.3). Furthermore, both H(.) and G(.) could be any model as long as the inputs and outputs match.
B.2 Constraint Generator
We design the following three tricks in the prompts to regulate and enhance the generation results. (1) Generation scope constraints: For space groups, we explicitly specify in the ”Instruction” part that valid space group types should range from 1 to 230; For Wyckoff positions, we provide candidate Wyckoff letters in the ”Input” part. (2) Numerical standardization: To minimize the impact of text length on LLMs, we standardized all input material properties by retaining three decimal places and adaptively determining appropriate units based on property types. (3) Text optimization: Instead of directly generating target space group type and Wyckoff positions, we incorporate the phrase ”the most suitable…” into the prompts, which has been empirically shown to improve generation accuracy.
B.3 Crystal Structure Generator
The crystal structure generator is designed to receive symmetry constraints, elemental composition, and target properties generated by constraint generator, subsequently generating desired crystal structures. We employ a diffusion model to jointly generate the lattice matrix $\bm{L}$ and fractional coordinates $\bm{F}$ under these conditions. Detailed descriptions of the forward diffusion process and reverse generation procedure are provided in Appendix C.1.
The overall architecture of our denoising model is illustrated in Figure 7, which derives two denoising terms $\hat{\epsilon}_{k}$ , $\hat{\epsilon}_{{F}^{{}^{\prime}}}$ under the constraints of elemental composition, crystal symmetry, and target properties. For feature embedding, we introduce continuous material property embeddings into input feature representations. Continuous properties, following Con-CDVAE [ye2024cdvae], are expanded by a Gaussian Radial Basis Function (RBF) and embedded by an MLP:
For the embedding of noised lattices and noised coordinates, we adopt the approach described in DiffCSP++ [jiao2024space]. Subsequently, we employ a multi-level cascade decoder to generate denoising terms. To enhance the model’s fitting capacity, we incorporate a multi-head cross-attention mechanism within the decoder architecture to fuse information from both the noised feature and the input feature. The structure of the decoder in the n-th layer is designed as
Appendix C Implementation Details
C.1 Hyper-parameters and Training Details
@@
We train CDVAE [xie2021crystal], DiffCSP++ [jiao2024space], and our crystal structure generator on three public datasets to validate the effectiveness of the structure generator. Using the dataset described in Section 2.2, we train CDVAE [xie2021crystal], DiffCSP++ [jiao2024space], FlowMM [miller2024flowmm], SymmCD [levy2025symmcd], Con-CDVAE [ye2024cdvae], MatterGen [zeni2023mattergen], and our CrystalGF to verify the validity of the overall framework. We split our data into 80% training, 10% validation, and 10% test sets while maintaining strict correspondence between space group and Wyckoff positions entries, this opeartion prevents data leakage by ensuring identical crystals are allocated to the same subset. To the best of our knowledge, Con-CDVAE and MatterGen represent the advanced multi-conditional generative models, which utilize chemical formulas with material properties and elemental composition with material properties as model inputs, respectively. Therefore, we conduct comparative evaluations of both models against our CrystalGF.
@@
For CDVAE, we apply the transformer-enhanced [vaswani2017attention] version of CDVAE as introduced in DiffCSP [jiao2024crystal], while maintaining consistent training parameters. For DiffCSP++, we train the denoising model with 6 layers, 512-dimension hidden layer, and 128-dimension Fourier embeddings and the training epochs are set to 3000, 1500, 1000 for Perov-5, MP-20, and MPTS-52. For FlowMM and SymmCD, we trained two models in unconditional mode and ab initio generative mode under default Settings, respectively. For MatterGen, we fine-tuning the pre-trained model with chemical system and property, including band gap and formation energy. The dimension of the property embedding and the chemical system multi-hot embedding is set to 512, and model is trained for 1000 epochs with an Adam optimizer with initial learning rate $5\times 10^{-6}$ and a Plateau scheduler with a decaying factor 0.6 and a patience of 100 epochs. For Con-CDVAE, we train model with chemical formula and property, including band gap and formation energy. The formula embedding module has 3 layers, 92-dimension input layer and 64-dimension hidden layer, the property embedding module has 3 layers and 64-dimension hidden layer and the model is trained for 1000 epochs with an Adam optimizer with initial learning rate $1\times 10^{-3}$ and a Plateau scheduler with a decaying factor 0.6 and a patience of 30 epochs.
@@
For our symmetry generator, We fine-tune four LLMs ,including Qwen2.5-7B-Instruct, Llama-3.1-8B-Instruct, DeepSeek-R1-8B, and GLM-4-9B-chat, using LoRA from Llamafactory [zheng2024llamafactory] framework. The LoRA fine-tuning specifically targets the qkv layers in transformer modules with 20 epochs and a $1\times 10^{-4}$ initial learning rate, the learning rate scheduler type is set to cosine annealing strategy with 0.1 warm up ratio. And the fine-tuning precision is set to FP32 throughout all computations.
@@
C.2 Symmetry Accuracy Details
@@
When constructing the confusion matrix for Wyckoff positions generation, we first identify matches where both the atomic order and generated letters coincide. Second, we distinguish cases where generated letters match despite atomic order discrepancies. Finally, atoms with inconsistent generation quantities are handled by padding with an ”other” category.
C.3 DFT Calculation Details
@@
CDVAE
Con-CDVAE
DiffCSP++
MatterGen
Constraint Generator
Structure Generator
NVIDIA A40 GPU
12GB
24GB
24GB
45GB
26GB + 30GB
30GB
10h
12h
30h
50h
9h + 18h
46h
C.4 Experiments Compute Resources
We present the computational resources consumed for training in Table 8, including GPU type, memory usage, and training duration. The resource consumption for CDVAE and DiffCSP++ corresponds to training with the MP-20 dataset mentioned in Appendix 2.3, while Con-CDVAE, MatterGen, the Constraint Generator, and the Structure Generator reflect the resource usage when trained on the dataset described in Appendix 2.2.
@@ C.4 Experiments Compute Resources
Appendix D Visualization of Generated Structures
@@
Appendix E Societal Impacts